OpenStudy (anonymous):

Solve this system of equations using any method. 4x – y = 5 x + y = 10 Answer choices (–1, –9) (3, 7) (5, 5) (4, 6)

OpenStudy (anonymous):

@dolphin1998 @imqwerty

OpenStudy (anonymous):

what do you get if you add the equations?

OpenStudy (anonymous):

without using the x and all tht just the numbers? @pgpilot326

OpenStudy (anonymous):

no, including the x and y... left + left = right + right. try that and see what results.

OpenStudy (anonymous):

@pgpilot326 that is confusing, explain

OpenStudy (anonymous):

what is left side of the first equation?

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

the whole left side

OpenStudy (anonymous):

4x-y

OpenStudy (anonymous):

now add to it the whole left side of the second equation

OpenStudy (anonymous):

x+y u can't

OpenStudy (anonymous):

@pgpilot326

OpenStudy (dolphin1998):

I think its (5,5)

OpenStudy (anonymous):

@dolphin1998 I was thinking that but it's confusing... :/

OpenStudy (anonymous):

you can combine like terms... please stop resisting and open your mind

OpenStudy (anonymous):

@pgpilot326 I am not resisting I am confused because I have no idea how to do Algebra work

OpenStudy (dolphin1998):

pgpilot326 she is confused. Please show some respect

OpenStudy (dolphin1998):

i would say its (5,5)

OpenStudy (anonymous):

@dolphin1998 thank you! that's what I'm thinking

OpenStudy (anonymous):

4x - y + (x + y) = 5 + 10 this simplifies to 4x + x -y + y = 15 and again 5x = 15 can you solve this?

OpenStudy (dolphin1998):

Your welcome

OpenStudy (dolphin1998):

No its (3,7)

OpenStudy (anonymous):

@dolphin1998 (3,7)

OpenStudy (dolphin1998):

Yay

OpenStudy (anonymous):

why do you only care about the answer and not how to solve? and @dolphin1998 how are you helping this person learn?

OpenStudy (anonymous):

@dolphin1998 she has done this before, but I am sure it is correct because it doesn't look as if any others fit

OpenStudy (dolphin1998):

She is my friend. Thanks for the concern. I am not a mean person so don't be mean to me. She probably knows how to she just doesn't get it

OpenStudy (anonymous):

i'm not being, just pointing out that what is going on here doesn't lead to learning.

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

@dolphin1998 exactly I am confused because I can't understand how we get these answers and all. @pgpilot326 ok

OpenStudy (dolphin1998):

ok

OpenStudy (anonymous):

@dolphin1998 I just thought up of how to find it. I can email my teacher mrs. hansen to see if she can help me understand it.

OpenStudy (dolphin1998):

That is a great Idea. Thats what I do

OpenStudy (anonymous):

@dolphin1998 I will let you know what happens next.

OpenStudy (dolphin1998):

okay good luck

OpenStudy (anonymous):

3x+y = 5 -------> y = 5-3x 4x - 7y = -10 ----------------- 4x - 7(5-3x) = -10 4x -35 + 21x = -10 4x + 21x = -10 + 35 25x = 25 x = 1 ------------> y = 5-3 = 2 https://answers.yahoo.com/question/index?qid=20100919160444AA6cYqO

OpenStudy (anonymous):

@Th3B3ast16 thank you!